OpenStudy (anonymous):

The sum of the ages of Brain,steve and phil is 53.brain's age now is twice steve's age five years ago,and brain is two years older than Phil.What is the age of the oldest? Please show me the steps. A.19 B.25 C.20 D.18 E.None of these

OpenStudy (anonymous):

its C

OpenStudy (anonymous):

to solve this analytically you can form 3 equations: x + y + z = 53 x = 2(y - 5) x = z + 2 where x = brain age, y = steves and z = phils age

OpenStudy (tommo_lcfc):

Equate the second and third equation as they have the same subject and substitute the second equation into the first one.

OpenStudy (anonymous):

x + y + z = 53 x - 2y = -10 subtract: 3y + z = 63..........(1) subtract x = z + 2 from the first equation y + 2z = 51..........(2) now you can solve equations (1) and (2) this gives y = 15 and x = 20 , z = x -2 = 18 oldest is 20

OpenStudy (tommo_lcfc):

Then you will get 2 simultaneous equations in y and z, namely 3y + z = 63 and 2y - z = 12. Eliminate z by adding the two together. 5y = 75 so y = 15. Use the second equation to find z so 2(15)-z = 12 so z = 18. Use equation 2 to find x. 2(15-5) = x = 2(10) = 20. Alternatively, just follow what @joeywhite did and you'll be fine! :)

OpenStudy (anonymous):

hope this helps

OpenStudy (anonymous):

Thanks